OpenStudy (anonymous):

The endpoints of one diagonal of a rhombus are (0, -8) and (8, -4). If the coordinates of the 3rd vertex are (1, 0), what are the coordinates of the 4th vertex?

OpenStudy (amistre64):

id say we can find the midpoint and use that to our advantage

OpenStudy (anonymous):

how would we do that?

OpenStudy (amistre64):

whats the middle between 0 and 8 ?

OpenStudy (anonymous):

-4?

OpenStudy (amistre64):

i dont believe there are any negative numbers between 0 and 8; but I would go with 4 :)

OpenStudy (amistre64):

whats the middle between -8 and -4?

OpenStudy (anonymous):

Oh haha you're right. not sure? -6?

OpenStudy (amistre64):

correct. ideally, we add the endpoints and divide it in half to get the mid point (0,-8) (8,-4) ------ (8,-12)/2 ; (4,-6)

OpenStudy (amistre64):

now we know the center of the thing, and that has to be in the center of the other vertexes right?

OpenStudy (anonymous):

Oh ok, and yeah

OpenStudy (amistre64):

(1,0) (x, y) ----- (x+1, y+0) /2 = (4,-6) therefore (x+1)/2 = 4 (y+0)/2 = -6 solve for x and y

OpenStudy (anonymous):

x=4 y=-6?

OpenStudy (amistre64):

um, not quite

OpenStudy (amistre64):

(x+1)/2 = 4 , *2 x+1 = 8 -1 -1 -------- x = 7 (y+0)/2 = -6 , *2 y+0 = -12 y = -12

OpenStudy (anonymous):

oh! so then x=-4?

OpenStudy (anonymous):

nvm haha 7!

OpenStudy (amistre64):

the magic 8ball says 7 lol

OpenStudy (anonymous):

haha so I saw. well I thank you for the help :)

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

Thank You